# Are Time-Indexed Foundation Models the Future of Time Series Imputation?

### v1

@@ Abstract
Foundation models for time series imputation remain largely unexplored. Recently, two such models, TabPFN-TS and MoTM, have emerged. These models share a common philosophy that places them within the family of time-indexed foundation models. This paper presents the first large-scale empirical study of these models for zero-shot imputation, which enables missing value recovery without retraining across a wide range of scenarios. We conduct extensive univariate experiments across 33 out-of-domain datasets ($\approx$ 1.3M imputation windows) and evaluate their ability to integrate covariates at inference time to improve accuracy without fine-tuning. Our results demonstrate that time-indexed foundation models are a powerful and practical step toward achieving general-purpose, zero-shot imputation for real-world time series.
00footnotetext: * Equal contribution
@@ 1 Introduction
Real-world time series from domains such as healthcare, industry, and climate science are often irregularly sampled or incomplete due to sensor failures and decentralized data collection (Schulz & Stattegger, 1997; Clark & Bjørnstad, 2004). Reliable imputation is thus a critical first step toward downstream tasks like forecasting, classification, or anomaly detection. Yet, while recent deep learning methods have advanced imputation performance (Cao et al., 2018; Du et al., 2023a), they typically lack robustness to distribution shifts and fail to generalize to out-of-domain data.
@@
A promising direction to overcome these challenges lies in continuous-time modeling, which learns a contextual representation $H(t)$ at each timestamp $t$, thereby casting imputation as a regression problem. Within this paradigm, time-indexed foundation models such as TabPFN-TS (Hoo et al., 2025) and MoTM (Le Naour et al., 2025) have recently been introduced, enabling zero-shot imputation through continuous-time representations. These models offer strong out-of-domain generalization without requiring retraining, making them particularly attractive for real-world applications where training data and computational resources are limited.
@@ 2.2 Time-Indexed Foundation Models for Imputation
MoTM (Le Naour et al., 2025) extends the continuous-time modeling paradigm by leveraging a sophisticated feature extraction mechanism inherited from the TimeFlow architecture (Le Naour et al., 2024). Its core principle is to represent any time series through a pre-trained basis of modulated Implicit Neural Representations (INRs).
@@ 3.1 Univariate Benchmark: Out-of-Domain Zero-Shot Imputation
3.1.1 Main Results
The aggregated results of our univariate benchmark are shown in Figure˜1, summarizing the mean normalized MAE across all 33 out-of-domain datasets. Each bar represents the overall imputation accuracy of a model averaged over the four missingness regimes. Models are grouped into three categories reflecting their underlying paradigm: (i) localmethods, which impute from within each series without any learned representation; (ii) task-specificmodels, trained in a supervised manner on each dataset; and (iii) foundation modelsevaluated in a fully zero-shot setting. To complement these aggregated results, we provide in Figure˜2 a critical difference (CD) diagram (Ismail Fawaz et al., 2019) summarizing the average ranks of all models across datasets and missingness scenarios.
@@ Results.
As shown in Figures˜1 and 2, three consistent trends emerge from the benchmark.
(i) Time-index foundation models lead the benchmark. Both TabPFN-TS and MoTM achieve the lowest aggregated normalized errors and the best average ranks, substantially outperforming all supervised and local baselines despite being evaluated in a fully zero-shot setting. The CD diagram further confirms that these two models are statistically superior compared to all others, except MoTM compared to SAITS. Among them, TabPFN-TS attains the top performance (MAE = 0.293, avg. rank = 1.35), suggesting that large-scale synthetic pretrained regressor combined with explicit temporal encodings confers a measurable advantage over the Ridge on top of learned continuous representations.
(ii) Task-specific models show limited robustness. While SAITS achieves competitive accuracy (0.386, avg. rank = 3.56), other supervised approaches such as BRITS, CSDI, or TimesNet lag behind, sometimes performing worse than simple local heuristics (e.g., Seasonal Naive, LOCF). This mixed behavior highlights the limited generalization capacity of fully supervised models, which tend to overfit dataset-specific temporal dynamics — particularly when training data is scarce.
(iii) Local baselines remain resilient. Classical approaches leveraging temporal priors still deliver reasonable performance in heterogeneous settings, as reflected in both their MAE and ranking consistency. However, the pronounced gap separating them from foundation models clearly illustrates the benefits of pretraining time-indexed models, which provide both accuracy and adaptability without retraining.
@@
Overall, the aggregated metrics and the rank-based analysis provide clear evidence that foundation models — particularly TabPFN-TS — deliver the most consistent and robust performance for zero-shot time-series imputation across diverse domains. MoTM also achieves an honorable level of accuracy despite relying on a comparatively simple regressor, underscoring the effectiveness of its pretrained time-indexed representations. Complete benchmark results are reported in Section˜C.1, while additional experiments on datasets with lower sampling rates (daily and weekly) are provided in Section˜C.4.
@@
Qualitative examples in Figure˜3 confirm these trends: both foundation models adjust their uncertainty bands (5–95 and 25–75 quantile ranges) to local signal variability. TabPFN-TS produces sharper, high-fidelity reconstructions that closely follow the ground truth, while MoTM yields smooth yet robust imputations with stable and well-calibrated uncertainty envelopes. Additional examples are provided in Section˜C.3.3.
@@ 3.2 Integration of Covariates in Zero-Shot OoD Imputation
For a model to be practically deployable in real-world applications, it must be capable of incorporating covariates, which often enhance predictive accuracy by providing additional contextual information. This requirement is recognized as one of the major challenges for foundation models in time series forecasting as well (Auer et al., 2025a). In this section, we investigate how TabPFN-TS and MoTM handle covariates in a zero-shot setting. We conduct experiments on three datasets where the dependence of the time series values on covariates ranges from weak to critical, in order to evaluate how efficiently they integrate this additional information into their imputations.
@@
The train / validation / test splits follow the same protocol as in Section˜3.1. A key difference is that at inference, the covariate is always assumed to be fully observed. We compare the following methods: TabPFN-TS and MoTM both with or without covariates, a ridge regression using only the covariate to predict the target variable, and SAITS. Results are showcased in Table˜2.
@@ Results.
As illustrated in Figures 4 and 5, incorporating covariates visually improves the reconstructions for both TabPFN-TS and MoTM. The covariate-enhanced versions capture missing intervals more smoothly and better follow short-term temporal variations within the four one-day gaps. We observe that the interpolations of TabPFN-TS align more closely with the ground truth and preserve sharper transitions across missing regions. In contrast, if MoTM consistently benefits from covariate information, its reconstructions sometimes struggle to capture sudden changes with great accuracy. More covariate plots are showcased in Appendix˜D, confirming that the most visible gains are observed on the Wind-France and PV-France datasets, while the impact remains limited for Load-France.
@@
As shown in Figure˜6, TabPFN-TS consistently achieves the lowest MAE but at the cost of substantially higher inference time, especially on large datasets. In contrast, MoTM offers a favorable balance between accuracy and efficiency, being up to two orders of magnitude faster while maintaining competitive performance. SAITS attains moderate error levels but requires significant computational overhead due to its supervised training phase. Finally, Linear is the fastest overall but shows markedly poorer accuracy, particularly on more complex datasets. These results highlight the trade-off between zero-shot generalization and computational cost, emphasizing MoTM as a scalable alternative to TabPFN-TS.
@@ A.2 Supervised imputers
All models implementations are taken from the PyPOTS (Du et al., 2023b) Python toolbox and trained on fixed-length windows where a subset of observed entries is randomly masked; each model reconstructs these masks from the remaining context (values ++ binary observation mask). Inputs are z-score normalized per variable, training minimizes the mean absolute error (MAE) on masked positions only, model selection uses validation MSE with early stopping, and test metrics are computed on the 4 missing points scenario detailed in Section 3.1. Hyperparameter choices were inspired by the default settings recommended by each method’s authors in their original papers and/or public implementations. Note that we use the Adam optimizer (Kingma & Ba, 2015) for all models with a batch size of 64, train for at most 50 epochs, and apply early stopping with a patience of 5 epochs.
@@ A.3 Time-Index Foundation models
In this section, we provide further architectural and implementation details for the two time-index foundation models evaluated in our experiments: MoTM and TabPFN-TS.
@@
MoTM (Mixture of TimeFlow Models) extends the continuous-time modeling paradigm by leveraging a sophisticated feature extraction mechanism inherited from the TimeFlow architecture (Le Naour et al., 2024). Its core principle is to represent any time series through a pre-trained basis of modulated Implicit Neural Representations (INRs).
Representation Learning via Modulated INR Basis.
Specifically, MoTM does not learn a single function for the time series but rather a basis of $K$ distinct INRs. Each INR is a small neural network, parameterized by a hypernetwork (Dupont et al., 2022), that maps a continuous time coordinate $t$ to a feature vector. These basis functions are "modulated" in the sense that their parameters are dynamically generated for each new segment, allowing them to capture a wide range of temporal patterns (e.g., trends, seasonalities, high-frequency oscillations) without being restricted to predefined frequencies like Fourier features. For any given timestamp $t$, the rich contextual representation $H(t)$ is formed by concatenating the outputs of all $K$ basis INRs evaluated at that time.
In-Context Imputation via Local Regression.
The key mechanism of MoTM for imputation lies in its local, in-context fitting procedure. Given a time series with missing values, MoTM considers a context window of observed points. It then fits a simple ridge regressor to learn the linear mapping from the high-dimensional representations $H(t)$ of these observed points to their actual values $x(t)$. This local regressor, fitted specifically on the available context, is then used to predict the values at any missing timestamp $t_{\text{miss}}$ by applying it to the corresponding representation $H(t_{\text{miss}})$.
This framework naturally extends to:
  1. 1.
    Integrate covariates with no retraining. Assuming full observation, additional contextual information available at timestamp $t$ are simply stacked to the target contextual representation $H(t)$. Ridge adaptation proceeds then in the same way as in the univariate setting.
  2. 2.
    Uncertainty quantification. By replacing the ridge regressor with a quantile regressor, MoTM can generate quantile predictions to construct confidence intervals around the imputed values.
In essence, MoTM decouples the problem: (i) a powerful, pre-trained feature extractor (the INR basis) captures complex temporal dynamics, (ii) while a simple, locally-fitted linear model performs the final task-specific regression. This design allows the model to adapt to the specific structure of each new time series without requiring any fine-tuning of its core components.
@@
TabPFN-TS Hoo et al. (2025) applies the continuous-time modeling philosophy by reframing the time series imputation task as a standard tabular regression problem. This allows the direct application of the powerful, pre-trained TabPFN model (Hollmann et al., 2025) for zero-shot time series analysis. The model’s design philosophy is conceptually inverse to that of MoTM: it pairs a simple, handcrafted feature representation with a highly expressive regression model.
Handcrafted Temporal Representations.
Unlike MoTM’s learned representations, TabPFN-TS employs a straightforward feature engineering approach. The contextual representation $H(t)$ for each timestamp $t$ is constructed by combining the normalized time index itself with a set of pre-defined Fourier basis functions (i.e., sine and cosine pairs). These Fourier features are chosen to capture key seasonalities expected in the data (e.g., daily, weekly). This method results in a simple, fixed feature set that explicitly encodes temporal position and periodicity, serving as the input for the regression model.
Imputation via In-Context Learning with TabPFN.
The core expressive power of TabPFN-TS resides in its regressor, the TabPFN model. TabPFN is a large transformer-based architecture pre-trained on hundreds of millions of synthetically generated tabular regression tasks. Its defining characteristic is in-context learning: at inference time, it ingests a set of observed data points—pairs of temporal features and their corresponding values, $(H(t_{obs}),x(t_{obs}))$ as a single "prompt." The model processes this entire context within its attention layers to infer the underlying functional relationship between the features and the series values. It then applies this inferred function to predict the values $x(t_{\text{miss}})$ for the query features $H(t_{\text{miss}})$ of missing timestamps, all within a single forward pass and without any gradient-based fine-tuning.
This framework naturally supports: (i) Integratation of covariates with no retraining.Covariates available at a timestamp $t$ are simply concatenated to the temporal representation $H(t)$, expanding the feature set that TabPFN uses for its in-context learning. (ii) Uncertainty quantification.TabPFNinherently supports probabilistic predictions. By configuring it to predict specific quantiles instead of the mean, we can generate confidence intervals for the imputed values, as demonstrated in Figure˜8.
In essence, TabPFN-TS also decouples the problem, but with an inverted philosophy compared to MoTM: (1) a simple, handcrafted feature extractor creates temporal features, (2) while a highly expressive, pre-trained transformer model performs the complex in-context regression. This design leverages the generalization capability of a large pre-trained model to solve the imputation task.

### v2

@@ Abstract
Foundation models for time series imputation remain largely unexplored. Recently, two such models, TabPFN-TS and MoTM, have emerged. These models share a common philosophy that places them within the family of time-indexed foundation models. This paper presents the first large-scale empirical study of these models for zero-shot imputation, which enables missing value recovery without retraining across a wide range of scenarios. We conduct extensive univariate experiments across 33 out-of-domain datasets ($\approx$ 1.3M imputation windows) and evaluate their ability to integrate covariates at inference time to improve accuracy without fine-tuning. Our results demonstrate that time-indexed foundation models are a powerful and practical step toward achieving general-purpose, zero-shot imputation for real-world time series. Code is available at https://github.com/taharnbl/tsfm_imputation.
@@ 1 Introduction
11footnotetext: Equal contribution.
Real-world time series from domains such as healthcare, industry, and climate science are often irregularly sampled or incomplete due to sensor failures and decentralized data collection (Schulz & Stattegger, 1997; Clark & Bjørnstad, 2004). Reliable imputation is thus a critical first step toward downstream tasks like forecasting, classification, or anomaly detection. Yet, while recent deep learning methods have advanced imputation performance (Cao et al., 2018; Du et al., 2023a; Nie et al., 2024), they typically lack robustness to distribution shifts and fail to generalize to out-of-domain data.
@@
A promising direction to overcome these challenges lies in continuous-time modeling, which learns a contextual representation $H(t)$ at each time-step $t$, thereby casting imputation as a regression problem. This approach is often referred to as time-index modelling (Woo et al., 2023). Within this paradigm, time-indexed foundation models such as TabPFN-TS (Hoo et al., 2025) and MoTM (Le Naour et al., 2025) have recently been introduced, enabling zero-shot imputation through continuous-time representations. These models offer strong out-of-domain generalization without requiring retraining, making them particularly attractive for real-world applications where training data and computational resources are limited.
@@ 2.2 Time-Indexed Foundation Models for Imputation
MoTM (Le Naour et al., 2025) extends the continuous-time modeling paradigm by leveraging a sophisticated feature extraction mechanism inherited from the TimeFlow architecture (Le Naour et al., 2024). Its core principle is to represent any time series through a pre-trained basis of modulated Implicit Neural Representations (INRs; see also (Li et al., 2025)).
@@ 3.1 Univariate Benchmark: Out-of-Domain Zero-Shot Imputation
3.1.1 Quantitative univariate results
The aggregated results of our univariate benchmark are shown in Fig.˜1, summarizing the mean normalized MAE across all 33 OoD datasets. Each bar represents the overall imputation error of a model averaged over the four missingness regimes. Models are grouped into three categories reflecting their underlying paradigm: (i) localmethods, imputing without any learned representation; (ii) task-specificmodels, trained in a supervised manner on each dataset; and (iii) foundation modelsevaluated in a fully zero-shot setting. In addition, Fig.˜2 presents a critical difference (CD) diagram (Demšar, 2006) that compares models via their average ranks across datasets and missingness scenarios. See also Table LABEL:tab:full-results in Appendix˜C for the full per-dataset results and Section˜C.4 for additional experiments on datasets with lower sampling rates (daily and weekly).
@@ Results.
As shown in Figures˜1 and 2, three takeaways emerge from the benchmark.
(i) Time-index foundation models lead the benchmark. TabPFN-TS achieves the lowest mean normalized MAE, followed by MoTM; both outperform all supervised and local baselines despite being evaluated fully zero-shot. In terms of ranking, the CD diagram indicates that TabPFN-TS attains the best average rank and is statistically superior to all competitors (MAE = 0.293, avg. rank = 1.35). While MoTM achieves the second-lowest aggregate MAE (MAE = 0.371), it ranks third by average rank (avg. rank = 3.62) and is not significantly different from SAITS under the CD comparison. These results suggest that large-scale synthetic pretrained regressor combined with explicit temporal encodings confers a measurable advantage over the Ridge on top of learned continuous representations.
(ii) Task-specific models show limited robustness. While SAITS achieves competitive accuracy (MAE = 0.386, avg. rank = 3.56), other supervised approaches such as BRITS, CSDI, or TimesNet lag behind, sometimes performing worse than simple local heuristics (e.g., Seasonal Naive, LOCF). This mixed behavior highlights the limited generalization capacity of fully supervised models, which tend to overfit dataset-specific temporal dynamics — particularly when training data is scarce.
(iii) Local baselines remain resilient. Classical approaches leveraging temporal priors still deliver reasonable performance in heterogeneous settings, as reflected in both their MAE and ranking consistency. For instance, Linear Interpolation achieves an average MAE of 0.506, against e.g 0.664 for CSDI or 0.677 for TimesNet. However, the pronounced gap separating them from foundation models clearly illustrates the benefits of pretraining time-indexed models, which provide both accuracy and adaptability without retraining.
@@
Overall, the aggregated metrics and the rank-based analysis provide clear evidence that foundation models — particularly TabPFN-TS — deliver the most consistent and robust performance for zero-shot time-series imputation across diverse domains. MoTM also achieves an honorable level of accuracy despite relying on a comparatively simple regressor, underscoring the effectiveness of its pretrained time-indexed representations. Finally, an ablation study in Appendix D indicates that TabPFN-TS’s gains are highly sensitive to the choice of time-index features: the selected handcrafted encoding (normalized time index and daily/weekly Fourier features) consistently outperforms variants that remove periodicities or replace them with random frequencies.
@@
Qualitative examples in Fig.˜3 confirm these trends: both foundation models adjust their uncertainty bands (5–95 and 25–75 quantile ranges) to local signal variability. TabPFN-TS produces sharper, high-fidelity reconstructions that closely follow the ground truth, while MoTM yields smooth yet robust imputations with stable and well-calibrated uncertainty envelopes. Additional examples are provided in Section˜C.3.3.
@@ 3.2 Integration of Covariates in Zero-Shot OoD Imputation
For a model to be practically deployable in real-world applications, it must be capable of incorporating covariates, which often enhance predictive accuracy by providing additional contextual information. This requirement is recognized as one of the major challenges for foundation models in time series forecasting as well (Auer et al., 2025a). In this section, we investigate how TabPFN-TS and MoTM integrate additional covariates in a zero-shot imputation setting. We conduct experiments on three datasets where the dependence of the time series values on covariates ranges from weak to critical. This section assumes that the covariate is fully observed; the scenario of partially observed covariates is discussed in Section˜E.2.
@@
The train / validation / test splits follow the same protocol as in Section˜3.1. We compare the following methods: TabPFN-TS and MoTM both with or without covariates, a ridge regression using only the covariate to predict the target variable, and SAITS. Results are showcased in Table˜2.
@@ Results.
As illustrated in Figures 4 and 5, incorporating covariates visually improves the reconstructions for both TabPFN-TS and MoTM. The covariate-enhanced versions capture missing intervals more smoothly and better follow short-term temporal variations within the four one-day gaps. We observe that the interpolations of TabPFN-TS align more closely with the ground truth and preserve sharper transitions across missing regions. In contrast, if MoTM consistently benefits from covariate information, its reconstructions sometimes struggle to capture sudden changes with great accuracy. More covariate plots are showcased in Appendix˜E, confirming that the most visible gains are observed on the Wind-France and PV-France datasets, while the impact remains limited for Load-France.
@@
As shown in Figure˜6, TabPFN-TS consistently achieves the lowest MAE but suffers from a substantially higher inference time, especially on large datasets. In contrast, MoTM offers a favorable balance between accuracy and efficiency, being up to two orders of magnitude faster while maintaining competitive performance. SAITS attains moderate error levels but requires significant computational overhead due to its supervised training phase. Finally, Linear is the fastest overall but shows markedly poorer accuracy, particularly on more complex datasets. These results highlight the trade-off between zero-shot generalization and computational cost, emphasizing MoTM as a scalable alternative to TabPFN-TS.
Practical deployment considerations.
As stated above, a practical limitation of the best model, TabPFN-TS, is its inference cost (see Section˜4.1): on an NVIDIA H100 GPU, a forward pass over a 672-step chunk takes approximately one second. This cost can be readily amortized in batch offline imputation or low-frequency pipelines, but may be prohibitive for real-time use cases or deployments operating on thousands of concurrent time series. Accordingly, the clearest deployment recommendation is compute-driven: use TabPFN-TS when GPU acceleration is available, and prefer MoTM in resource-constrained settings or when throughput across many sequences becomes the primary bottleneck. A more detailed discussion is provided in Appendix D.
@@ 5 Conclusion and Discussion
Acknowledgements
We would like to thank the authors of all the time series datasets listed in Appendix˜B, as well as the authors of the PyPOTS Python toolkit (Du et al., 2023b), for enabling this work by their strong contribution to the open-source time series research community. Finally, we thank the reviewers for their thoughtful reviews that helped improve the paper.
@@ A.2 Supervised imputers
All models implementations are taken from the PyPOTS (Du et al., 2023b) Python toolbox and trained on fixed-length windows where a subset of observed entries is randomly masked; each model reconstructs these masks from the remaining context (values ++ binary observation mask). Inputs are z-score normalized per variable, training minimizes the mean absolute error (MAE) on masked positions only, model selection uses validation MSE with early stopping, and test metrics are computed on the 4 missing points scenarios detailed in Section 3.1. For all baselines, we use the default hyperparameter settings recommended by the original authors in their papers or public implementations. In addition, for the strongest supervised baseline, SAITS, we perform an extensive hyperparameter study on 11 representative datasets (see Appendix˜F). These experiments show that, while careful tuning can yield modest improvements, the default hyperparameter configuration already provides competitive performance at a substantially lower computational cost than tuning each model separately on every dataset. Note that we use the Adam optimizer (Kingma & Ba, 2015) for all models with a batch size of 64, train for at most 50 epochs, and apply early stopping with a patience of 5 epochs.
@@ A.3 Time-Index Foundation models
This section complements Section˜2.2 by providing further implementation details for the two time-index foundation models evaluated in our experiments: MoTM and TabPFN-TS.
@@ A.4 Excluded baselines
Our main univariate benchmark includes two foundation models, six deep supervised imputers and three local baselines. Although extensive, this selection could be complemented by several other models. Below we motivate why other relevant foundation models (MOMENT and NuwaTS), deep supervised models and local baselines (Spline Interpolation) where not included in the main experiments. For the sake of completeness, evaluation of these excluded baselines on a representative subset of the benchmark is carried out in Section˜A.4.1, confirming that they lack robustness to perform well on diverse datasets and settings.
@@
NuwaTS.
NuwaTS is another Transformer-based model specifically designed to address zero-shot time series imputation by repurposing pretrained language models (Cheng et al., 2024). Although the public codebase [^1] Code available at https://github.com/Chengyui/NuwaTS. does not provide any off-the-shelf pretrained models, the official demonstration [^2] Demo available at https://colab.research.google.com/drive/1jjM6g4N7AqyHjYawZWJdbFgNY7p4ZtGY?usp=sharing. allows experimenting with a version pretrained on fixed segments of length 96. To integrate NuwaTS into our evaluation pipeline, we adapted this demonstration model by splitting each considered time series into multiple 96-length segments from which NuwaTS could extract shared information. This workaround is clearly non-ideal but enables a preliminary zero-shot assessment of the model under comparable conditions. Consistently with the rest of our analysis, these experiments indicate that NuwaTS trails substantially behind time-indexed foundation models.
@@
Spline interpolation.
Cubic Spline interpolation is a simple non-linear imputation method based on piecewise polynomials of third-order. Although it produces smooth interpolations and an accurate fit, this method is known to be prone to overfitting under sparse or noisy observations.
A.4.1 Additional experiments
Setting.
We conduct supplementary experiments to evaluate MOMENT (Goswami et al., 2024), NuwaTS (Cheng et al., 2024) and a Cubic Spline Interpolation. MOMENT and NuwaTS operate in a fully zero-shot setting. They are compared against TabPFN-TS, MoTM and Linear Interpolation.Table˜3 shows their performance on eleven representative univariate datasets (see Appendix˜B) and four missingness scenarios (see Section˜3.1).
Table 3: Complete MAE results of additional baselines MOMENT, NuwaTS and CubicSpline on 11 datasets. For each setting, the best score (lowest) is in bold and the second best is underlined.
Dataset Setting
TabPFN-TS
MoTM
MOMENT
NuwaTS
CubicSpline
Linear
BDG2–Bear Pointwise 1 0.171 0.202 1.026 0.349 0.281 0.229
Pointwise 2 0.223 0.240 1.026 0.436 0.447 0.330
Blocks 1 0.272 0.332 1.075 0.519 3.033 0.857
Blocks 2 0.280 0.336 1.077 0.535 3.000 0.861
BDG2–Rat Pointwise 1 0.196 0.231 0.811 0.352 0.317 0.247
Pointwise 2 0.256 0.273 0.814 0.428 0.504 0.334
Blocks 1 0.349 0.400 0.808 0.516 4.099 0.743
Blocks 2 0.355 0.402 0.808 0.524 4.085 0.744
Covid19 Energy Pointwise 1 0.075 0.099 1.024 0.287 0.124 0.163
Pointwise 2 0.132 0.127 1.021 0.381 0.272 0.292
Blocks 1 0.202 0.222 1.075 0.457 1.689 0.949
Blocks 2 0.201 0.232 1.073 0.474 1.107 0.928
GFC12 Load Pointwise 1 0.143 0.189 0.985 0.348 0.202 0.232
Pointwise 2 0.237 0.231 0.985 0.436 0.419 0.370
Blocks 1 0.353 0.382 1.043 0.486 2.622 0.810
Blocks 2 0.363 0.387 1.047 0.499 2.682 0.814
Hog Pointwise 1 0.196 0.240 0.979 0.320 0.287 0.216
Pointwise 2 0.260 0.286 0.979 0.375 0.426 0.280
Blocks 1 0.396 0.458 1.029 0.491 3.043 0.555
Blocks 2 0.406 0.464 1.029 0.502 2.848 0.564
Jena Weather 10T Pointwise 1 0.086 0.190 0.982 0.143 0.101 0.080
Pointwise 2 0.101 0.207 0.981 0.165 0.132 0.096
Blocks 1 0.378 0.453 1.030 Nan 4.774 0.527
Blocks 2 0.384 0.459 1.024 Nan 4.246 0.544
Jena Weather 1H Pointwise 1 0.156 0.224 0.990 0.300 0.194 0.170
Pointwise 2 0.214 0.272 0.983 0.364 0.343 0.243
Blocks 1 0.366 0.463 1.078 0.475 2.698 0.552
Blocks 2 0.370 0.464 1.055 0.489 2.605 0.555
Oikolab Weather Pointwise 1 0.150 0.227 0.984 0.318 0.176 0.164
Pointwise 2 0.228 0.278 0.984 0.391 0.316 0.248
Blocks 1 0.449 0.529 1.043 0.556 1.732 0.638
Blocks 2 0.454 0.534 1.043 0.562 1.801 0.636
PDB Pointwise 1 0.062 0.094 0.984 0.269 0.120 0.191
Pointwise 2 0.119 0.121 0.977 0.339 0.278 0.338
Blocks 1 0.171 0.197 1.051 0.352 1.200 0.982
Blocks 2 0.177 0.201 1.053 0.382 1.801 1.006
Pedestrian Counts Pointwise 1 0.150 0.196 1.015 0.433 0.404 0.329
Pointwise 2 0.200 0.239 1.015 0.502 0.669 0.447
Blocks 1 0.172 0.254 1.070 0.498 3.644 1.001
Blocks 2 0.178 0.260 1.069 0.513 5.382 1.000
Weather Pointwise 1 0.257 0.298 0.966 0.360 0.359 0.259
Pointwise 2 0.311 0.342 0.965 0.411 0.521 0.322
Blocks 1 0.456 0.494 1.027 0.527 2.992 0.620
Blocks 2 0.455 0.488 1.018 0.528 2.998 0.618
Mean (Std) 0.252 (0.114) 0.300 (0.122) 1.002 (0.070) 0.419 (0.101) 1.613 (1.535) 0.502 (0.286)
Results.
Table˜3 shows the z-normalized MAE scores across all eleven datasets and four missingness settings, enlighting why MOMENT, NuwaTS and Cubic Spline Interpolation where excluded from the main benchmark. (i) MOMENTfails to produce accurate predictions across all datasets and settings, confirming that it is not a suitable zero-shot time-series imputation model. (ii) NuwaTSoutperforms MOMENT yet lags significantly behind TabPFN-TS on all datasets and settings and behind MoTM on ten out of the eleven datasets. Altogether, these observations strengthen the time-indexed approaches as strong alternatives to Transformer-based foundation models for time series imputation. (iii) Finally, the Cubic Spline Interpolation suffers from severe overfitting in challenging scenarios, such as high missingness rate (Pointwise 2, 70% of missing values) or block scenarios. By contrast, the Linear Interpolation provides more robust predictions and outperforms Cubic Splines in 39 out of 44 settings.